In was just wondering if I got this right.
The problem is 5x+5y=5
3y=-6x+12
The answer I got was x=1 y=-2 please help I don't know if I'm right out not
Systems of equations problem using addition/elimination

Found 2 solutions by mananth, MathTherapy:
Answer by mananth(16946)

5 x + 5 y = 5 .............1
Total value
6 x + 3 y = 12 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 5
-15 x -15 y = -15
30 x + 15 y = 60
Add the two equations
15 x = 45
/ 15
x = 3
plug value of x in (1)
5 x + 5 y = 5
15 + 5 y = 5
5 y = 5 -15
5 y = -10
y = -2

Answer by MathTherapy(10552)

In was just wondering if I got this right.
The problem is 5x+5y=5
3y=-6x+12
The answer I got was x=1 y=-2 please help I don't know if I'm right out not
Systems of equations problem using addition/elimination
To determine if your answers are correct, simply substitute the values for the variables into any of the original equations. If they satisfy (make true) the equation,
then the answers are correct. If you do that here, you'll see that your answers are INCORRECT.
You should do it like this:
1) Factor out GCF, 5 in the 1st equation to get: x + y = 1 ------- eq (i)
2) Factor out GCF, 3 in the 2nd equation to get: y = - 2x + 4____2x + y = 4 ------- eq (ii)
3) Subtract eq (i) from eq (ii) to eliminate y and find the value of x
4) Substitute the value of x into any of the 2 original equations to determine the value of y
5) Conduct a test as I'd mentioned earlier to determine if your values are correct